Two planes leave St. Louis, one lying east and the other lying west. The westbound plane travels 100 mph faster than the eastbound plane, and after 1.5 hours they are 750 miles apart. Find the speed of each plan.
Answer by lwsshak3(11628)

Two planes leave St. Louis, one flying east and the other flying west. The westbound plane travels 100 mph faster than the eastbound plane, and after 1.5 hours they are 750 miles apart. Find the speed of each plan.
let x=speed of eastbound plane
x+100=speed of westbound plane
distance=travel time*speed
1.5x+1.5(x+100)=750
1.5x+1.5x+150=750
3x=600
x=200
x+100=300
=speed of eastbound plane=200 mph
speed of westbound plane=300 mph